OpenStudy (anonymous):

Express the following using inverse notation 1/(5x^3)

OpenStudy (freckles):

what do you mean by inverse notation?

imqwerty (imqwerty):

do u mean like - 1/x = x^(-1)

OpenStudy (freckles):

\[y=\frac{1}{5x^3}\] I think you mean to write this maybe and you are looking for inverse function?

OpenStudy (freckles):

I have no clue what you mean really

OpenStudy (anonymous):

The problem says "Express the following in inverse notation"

OpenStudy (freckles):

You can please define that for me.

OpenStudy (anonymous):

ill take a picture of it

imqwerty (imqwerty):

i think you are talking about multiplicative inverse? u in 6or7th grade?

OpenStudy (anonymous):

Alright this is it

OpenStudy (anonymous):

No Ive just been out of school for five years

OpenStudy (freckles):

\[\frac{1}{a}=a^{-1}\] @imqwerty made a great guess earlier

OpenStudy (freckles):

oh I don't know why that is called inverse notation in your class you are just writing the expression with negative exponents instead *

imqwerty (imqwerty):

yea :D

OpenStudy (anonymous):

So where is a good place to practice this because I really don't have a clue how you go the answer.

OpenStudy (freckles):

you see the exponent 3

OpenStudy (freckles):

on the bottom there?

OpenStudy (anonymous):

yup

OpenStudy (freckles):

\[\frac{1}{5x^3}=\frac{1}{5} \frac{1}{x^3}=\frac{1}{5} x^{-3} \text{ use the rule I gave }\]

OpenStudy (freckles):

bring that little part to the numerator and change the sign of the exponent

OpenStudy (anonymous):

That makes sense now.

OpenStudy (freckles):

another example: \[\frac{5}{3(x+1)} \\ \frac{5}{3} \frac{1}{x+1}=\frac{5}{3} \frac{1}{(x+1)^1} \\ =\frac{5}{3}(x+1)^{-1}\]

OpenStudy (anonymous):

ah I see. This was easier than what I made it seem.

OpenStudy (rushwr):

\[\frac{ 1 }{ 5x ^{3} } = y\] \[5y= \frac{ 1 }{ x ^{3} }\]

OpenStudy (rushwr):

\[\sqrt[3]{5y} = x\]

imqwerty (imqwerty):

hey is it like the minus sign of the exponent is not written or m getting something wrng.....this has happened to me before http://prntscr.com/8bpkrt

OpenStudy (rushwr):

I have this link for you if u aren't sure of what you are doing ! https://www.mathsisfun.com/sets/function-inverse.html

OpenStudy (anonymous):